Solve the equation by making an appropriate substitution. 14) (x2 - 2x) - 23x2 - 2x)+ 120 - 0 answer {-3, -2, 5,4} Solve theshow work for both
Answer #1

And : 14). (x²-2x) * - 23 (x - 2x) +120=0 Let x2-2x= u 5o u² - 23 ut 1200 u²-84-15 u + 1200 u(4-8) -15 (4- 8) = 0 50 (4-15) ((a) when u= 15 22 - 2x - 15-0 2- 52+ 3x-15-0 (2-5) +3(2-5)=0 2=5, - -3 (6) when u= 8 22-22-8-0 x² 40+2x-850 (2-4) + 2(x-4)=0